OpenStudy (anonymous):

Will medal What is the period of the sinusoid given by y= -2sin ((2pie/7)x)

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

2pi/7 times what = 2pi

OpenStudy (danjs):

\[\frac{ 2 \pi }{ 7 } * x = 2 \pi\]

OpenStudy (danjs):

What is x, that is your period

OpenStudy (anonymous):

7

OpenStudy (danjs):

yep, that is all you have to do

OpenStudy (anonymous):

Ok thank you

OpenStudy (danjs):

the period of a normal sin(x) is 2pi,

OpenStudy (anonymous):

7